OpenStudy (anonymous):

Find the inverse of f(x)=2x^3-1

OpenStudy (campbell_st):

start by swapping x and y so it becomes \[x = 2y^3 - 1\] noe make y the subject... that will be the inverse

OpenStudy (anonymous):

In order to find the inverse of a function, first, replace f(x) with y. Next, change your y into an x, and your x into a y. Now solve for y. Example: \[f(x) = 3x-4\]\[y=3x-4\]\[x=3y-4\]\[x+4=3y\]\[\frac{x+4}{3}=y\]\[f^{-1}(x)=\frac{x+4}{3}\]

OpenStudy (anonymous):

I already know those steps. I'm just confused in how to work out diving 2y^3 since it's cubed

OpenStudy (campbell_st):

ok... start by adding 1 to both sides \[x + 1 = 2y^3\] now divide by 2 \[\frac{x + 1}{2} = y^3\] lastly take the cube root of both sides hope it helps